# EW corrections and Heavy Boson Radiation at a high-energy muon collider

### v1

@@ Abstract
TIF-UNIMI-2024-14, COMETA-2024-23
@@ 1 Introduction
In recent years a novel interest for a muon collider has arisen, motivated by its great potential for the investigation of the fundamental interactions of Nature MuonCollider:2022xlm ; Aime:2022flm ; Black:2022cth ; Maltoni:2022bqs ; Belloni:2022due ; Accettura:2023ked . A key aspect of a muon collider is the possibility of accelerating elementary particles at energies of several TeV’s Delahaye:2019omf ; Bartosik:2020xwr ; Schulte:2021hgo ; Long:2020wfp ; MuonCollider:2022nsa ; MuonCollider:2022ded ; MuonCollider:2022glg , leading to the possibility to probe fundamental interactions at unprecedented energies. It offers the potential to significantly advance our understanding of fundamental particle physics, enabling in-depth studies of the Standard Model (SM) Higgs boson Han:2020pif ; Buttazzo:2020uzc ; Han:2021lnp ; Reuter:2022zuv ; Celada:2023oji ; Forslund:2022xjq ; Chen:2022yiu ; Ruhdorfer:2023uea ; Han:2023njx ; Dermisek:2023rvv ; Liu:2023yrb ; Li:2024joa , searches for beyond-the-SM (BSM) heavy Higgs bosons Han:2022edd ; Bandyopadhyay:2020otm ; Han:2021udl ; Bandyopadhyay:2024plc ; Ouazghour:2023plc ; Jueid:2023qcf , investigations of dark matter Han:2020uak ; Capdevilla:2021fmj ; Han:2022ubw ; Cesarotti:2024rbh ; Asadi:2023csb ; Belfkir:2023vpo ; Jueid:2023zxx , constraints on lepton-universality violation Huang:2021nkl ; Huang:2021biu ; He:2024dwh , exploration of the muon $g-2$ anomaly Buttazzo:2020ibd ; Capdevilla:2020qel ; Yin:2020afe ; Capdevilla:2021rwo , and tests of a wide range of new physics scenarios Gu:2020ldn ; Costantini:2020stv ; Bandyopadhyay:2021pld ; Qian:2021ihf ; Bao:2022onq ; Ghosh:2023xbj ; Dermisek:2023tgq ; Altmannshofer:2023uci ; Liu:2023jta ; Sun:2023cuf ; Kwok:2023dck ; Cesarotti:2023sje ; Ema:2023buz ; Li:2023tbx .
@@
A muon collider is therefore both a discovery and precision machine. In particular, precision physics relies on both a clean experimental environment (no QCD in the initial state) and precise and reliable theoretical predictions. Thus, as has been the case for LEP, the calculation of electroweak (EW) corrections will be paramount for theoretical predictions in muon collisions. In fact, unlike other possible future electron-positron colliders, e.g. the Future Circular Collider (FCC-ee) FCC:2018evy ; Bernardi:2022hny and the Circular Electron-Positron Collider (CEPC) CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi ; An:2018dwb ; CEPCAcceleratorStudyGroup:2019myu ; CEPCPhysicsStudyGroup:2022uwl ; CEPCStudyGroup:2023quu , at a high-energy muon collider the scope and relevance of EW corrections will be much broader, especially for a 10 TeV (or higher) collision machine. EW corrections at high energies can be very large, even of ${\cal O}(1)$ w.r.t. the leading-order (LO) prediction, and therefore are expected to be unavoidable in any phenomenological study, not only those regarding precision. The origin of such enhancements is the so-called EW Sudakov logarithms (EWSL), which involve logarithms of ratios of the form $Q/M_{W}$, where $Q$ is any of the scale of the process, like the energy of the collider $\sqrt{S}$, and $M_{W}$ is the $W$ boson mass.
@@
In this work, we precisely investigate this issue for the case of a muon collider at 3 and 10 TeV. We focus on direct production processes (also denoted in the literature as muon-muon annihilation), $\mu^{+}\mu^{-}\to F$, where the invariant mass of the final state is close to the energy of the collider and therefore VBF configurations are suppressed. In these configurations the only relevant PDFs are the ones of the (anti)muon in the (anti)muon, which accounts for effects from QED initial-state radiation (ISR). Moreover, our focus is not on precision physics but on large effects, as those expected from EWSL at high energies. We exploit the SM as a “test case”, but our conclusions are instructive also for a general BSM scenario. [^2] The case of SMEFT will be addressed in detail in an upcoming publication elfaham:xxx . All our calculations are performed via the MadGraph5_aMC@NLO framework Alwall:2014hca ; Frederix:2018nkq ; Pagani:2021vyk ; Bertone:2022ktl , in a completely automated approach.
@@
Then we approximate the resummation of EWSL via a simple exponentiation and, after additively matching them to the exact NLO EW, we compare this prediction with the exact NLO EW itself. We investigate when resummation is expected to be relevant only for precision studies and when instead is unavoidable in order to obtain sensible predictions, especially positive defined.
@@ 2.1 The automation of NLO EW corrections
Given a physical observable, typically a cross section, the so-called NLO QCD and NLO EW corrections correspond to the exact ${\cal O}(\alpha_{s})$ and ${\cal O}(\alpha)$ corrections, respectively, to its LO prediction.  [^3] For the processes that we will consider in this paper, a single coupling combination contributes at LO. See Footnote 1 for the more general case. Such corrections involve the calculation of one-loop amplitudes, their renormalisation, the regularisation of IR divergencies, and the combination of virtual as well as real-emission contributions in order to cancel them.
@@ 3.1 Calculation set-up
The two classes of processes entail completely different kinematics, especially in the bulk of the associated cross sections. Direct production is dominated by the phase-space region $s\simeq S$, where $\sqrt{s}$ is the total energy of the partonic process in its rest frame while $\sqrt{S}$ is the collider energy. Configurations with $s\neq S$ are induced by the emissions of photons, in particular the initial-state-radiation (ISR), which is taken into account in the collinear limit directly via PDF evolution of the (anti)muon in the (anti)muon $\Gamma_{\mu^{\pm}/\mu^{\pm}}$ or otherwise via the NLO EW corrections. VBF production is instead dominated by the phase-space region $\sqrt{s}\simeq m(F)\simeq\sum_{i\in F}m_{i}$, where $i$ is any particle that is part of the final state $F$. In other words, The $\mu\mu^{-}/\nu_{\mu}\bar{\nu}_{\mu}$ additional pair in the final state carries away most of the energy of the colliding muons and it typically does it along the beam pipe axis. The hard process is in fact $VV\rightarrow F$, and the leading contributions, especially at very high energies, can be simulated in the so-called Effective-Vector-boson-Approximation (EVA) Dawson:1984gx ; Kane:1984bb , which has already been implemented in MadGraph5_aMC@NLO Ruiz:2021tdt . Instead of simulating $22\rightarrow n+$ 2 processes, where $n$ is the multiplicity of $F$, $2\rightarrow n$ matrix elements are sufficient ($VV\rightarrow F$) and the “$V$ in the muon” can be modeled similarly to what is done in the case of the photon in the lepton in the Weizsäcker–Williams approximation vonWeizsacker:1934nji ; Williams:1934ad . Not only, as already mentioned Sec. 2.1, these effects can also be resummed and taken into account in a PDF formalism, as shown in Refs. Han:2020uid ; Han:2021kes ; Garosi:2023bvq . [^8] It is interesting to note that since, at least for low multiplicities, $m(F)\simeq\sum_{i\in F}m_{i}\simeq M_{W}$, not only logarithms of the form $\log(s/M_{W}^{2})$ entering the PDFs are important but also power corrections of the form $(m(F)/M_{W})^{n}$ to the matrix elements cannot be neglected. However, this aspect is beyond the scope of this paper and it will not be investigated here.
@@
In this work we want to study the phenomenology of EW corrections at muon colliders for hard scattering processes at high energies, i.e., for the direct production mechanism.  [^9] The calculation of EW corrections for the direct production of multi-boson final states has been performed also in Ref. Bredt:2022dmm Unless a very high final-state multiplicity is chosen, in the SM for both 3 and 10 TeV collisions the two scales $m(F)$ and $S$ are very well separated, with $m(F)\ll S$. Thus, also when NLO EW corrections are taken into account, the two classes of processes (direct production and VBF) can be studied independently. In order to do so, unless differently specified, in all the results of the paper we apply the following cut on the final state $F$:
@@
In order to also be as close as possible to a realistic experimental set-up, since we consider only undecayed particles that are typically boosted, we require that they are both separated in angle among them and also with the beam pipe axis. In particular, for any particle $X$ and $Y$ that are part of the final state $F$ we require that
@@
where cuts are applied after the recombination of photons described before.
@@ 3.2.2 EWSL and their approximate resummation
The r.h.s. of Eq. (3.20) says that if $\sigma_{\rm EXP_{EW}}$ is expanded in powers of $\alpha$ the NLO EW prediction is captured exactly, while beyond ${\cal O}(\alpha)$ the resummed tower of EWSL of order $\alpha^{n}\log^{k}(s/M_{W}^{2})$ with $n>1$ and $k=2n,2n-1$ is approximated via simple exponentiation. We stress again that we do not claim we are doing NLL resummation of EWSL. We instead want to study when and if it is necessary this procedure, by comparing $\delta_{\rm NLO_{EW}}$ with the relative corrections induced by $\sigma_{\rm EXP_{EW}}$, namely
@@ 3.3 List of aspects investigated in this work
    When NLO EW corrections reach, or even worse surpass, the relative size of $-100\%$ of the LO, resummation of EWSL does not concern precision but the sensibility of the predictions. We want to investigate when we should expect such situations at 3 and 10 TeV collisions.
@@ 4 Numerical results
In this section we will also start to describe features of the EW corrections that are distinctive for direct production processes at high-energy lepton colliders and quite different from the case of a high-energy hadronic machine such as LHC or FCC-hh (see, e.g., Refs. Mangano:2016jyj ; Azzi:2019yne for overviews of EW corrections at high-energy hadronic machines). We remind the reader that all calculations have been performed for SM processes, which are the only ones that can be calculated at $\rm NLO_{EW}$ accuracy with automated tools. However, many of the conclusions from this study, especially those stating limitations of some approaches, can be clearly generalised also to BSM scenarios. The SM has to be regarded only as a test case in this respect.
@@ 4.1.1 Importance of ${\rm SDK_{weak}}$ versus ${\rm SDK_{0}}$
Using the same colour code of Ref. Pagani:2021vyk , in the main panel we show LO (blue), $\rm NLO_{EW}$ (orange), ${\rm SDK_{weak}}$ (red), and ${\rm SDK_{0}}$ (green) predictions. [^15] The rigorous definition of these quantities can be found in Sec. 3.2. In the first inset, we plot the relative corrections induced by such approximations w.r.t. the LO predictions, i.e., the quantities $\delta_{X}$ defined in Eqs. (3.10), (3.14) and more in general (3.26). In such inset it is possible to appreciate the size and the sign of EW corrections, either calculated at exact $\rm NLO_{EW}$ accuracy or via Sudakov approximation(s). Then, in the second inset, we plot the difference $\delta_{X}-\delta_{\rm NLO_{EW}}$ for the two cases $X={\rm SDK_{weak}}$ and $X={\rm SDK_{0}}$ in order to test their accuracy. Rather than the minimisation of this quantity, the validity of the Sudakov approximation consists in having a small constant difference ($|\delta_{X}-\delta_{\rm NLO_{EW}}|\simeq\alpha$) over the full spectrum, i.e. a horizontal line, see also Eqs. (3.18) and (3.19) and text around it. Indeed, ${\cal O}(\alpha)$ contributions are expected to be present, while non-horizontal lines indicate an (at least) logarithmic-enhanced contribution that is not captured. Such contribution may accidentally compensate the ${\cal O}(\alpha)$ constant term and lead to $\delta_{X}-\delta_{\rm NLO_{EW}}\simeq 0$ for particular phase-space regions, but this is not to be regarded as an indication of the validity of the approximation. That said, a large constant difference ($|\delta_{X}-\delta_{\rm NLO_{EW}}|\gg\alpha$), however, also points to logarithms that are not correctly captured. In particular, at a high-energy lepton collider, the direct-production processes studied in this work and characterised by $\sqrt{s}\simeq\sqrt{S}$ may show such effects induced by missing large double logarithms (see Eq. (2.3)) of the form $L(s,M_{W}^{2})$, [^16] It is easy to see in Sec. 4.1 of Ref. Pagani:2021vyk that, at variance with the ${\rm SDK_{0}}$ scheme, the ${\rm SDK_{weak}}$ scheme implies the substitution $C_{\rm EW}\to C_{\rm EW}-Q^{2}$, where $Q$ is the charge of the particle considered, in the prefactor of the double-logarithm of the form $L(s,M_{W}^{2})$ entering the formula of $\delta_{\rm DL}$ in Eq. (3.14). Similar effects are also present in the single logarithms and lead to large constants when a fixed $s$ is considered. which are therefore large constants for the full phase-space.
@@
Figure 5: Same as Fig. 1, for the Higgs boson $p_{T}$ distribution in $\mu^{+}\mu^{-}\to ZH$.
Finally, in Fig. 5 we show the transverse-momentum distribution for the Higgs boson in $\mu^{+}\mu^{-}\to ZH$ production, $p_{T}(H)$. Similarly to the case of $t\bar{t}$ production in Fig. 1, we observe a much less flat LO prediction w.r.t. the case of $WW$ and $ZZ$ production. As in the case of $t\bar{t}$ production, and unlike $WW$ and $ZZ$, the tree-level amplitude features only an $s$-channel diagram. As in $t\bar{t}$ production, and even more, $\delta_{\rm NLO_{EW}}$ is quite flat over the full spectrum. We also observe that the difference $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$ is larger, of the order of $10\%$ both for 3 and 10 TeV collisions. Instead, in the case of $\delta_{{\rm SDK_{0}}}-\delta_{\rm NLO_{EW}}$, such difference is of the order of $15\%$ at 3 TeV and $20\%$ at 10 TeV. Thus, the accuracy of $\delta_{{\rm SDK_{0}}}$ is not only worse but also energy-dependent. To the best of our understanding, the $\mu^{+}\mu^{-}\to ZH$ has large non-logarithmic-enhanced contributions at ${\cal O}(\alpha)$ and on top of that the $\delta_{{\rm SDK_{0}}}$ wrongly approximates the double logarithms of the form $L(s,M_{W}^{2})\simeq L(S,M_{W}^{2})$ from the photon exchange among the muons in the initial state. Notice that the difference $\delta_{{\rm SDK_{weak}}}-\delta_{{\rm SDK_{0}}}$ is the same observed, both at 3 and 10 TeV, for $\mu^{+}\mu^{-}\to ZZ$ production in Fig. 4, for which the same argument was presented.
We have also calculated and analysed the $m(ZH)$ distribution. The situation is similar to the one observed for the $m(ZZ)$ distribution, but the $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$ remains constant at the order of $10\%$ as discussed for the $p_{T}(H)$ distribution.

### v2

@@ Abstract
TIF-UNIMI-2024-14, COMETA-2024-23, IRMP-CP3-25-08
arxiv: 2409.09129
@@ 1 Introduction
In recent years a novel interest for a muon collider has arisen, motivated by its great potential for the investigation of the fundamental interactions of Nature MuonCollider:2022xlm ; Aime:2022flm ; Black:2022cth ; Maltoni:2022bqs ; Belloni:2022due ; Accettura:2023ked . A key aspect of a muon collider is the possibility of accelerating elementary particles at energies of several TeV’s Delahaye:2019omf ; Bartosik:2020xwr ; Schulte:2021hgo ; Long:2020wfp ; MuonCollider:2022nsa ; MuonCollider:2022ded ; MuonCollider:2022glg , leading to the possibility to probe fundamental interactions at unprecedented energies. It offers the potential to significantly advance our understanding of fundamental particle physics, enabling in-depth studies of the Standard Model (SM) Higgs boson Han:2020pif ; Buttazzo:2020uzc ; Han:2021lnp ; Reuter:2022zuv ; Celada:2023oji ; Forslund:2022xjq ; Chen:2022yiu ; Ruhdorfer:2023uea ; Han:2023njx ; Dermisek:2023rvv ; Liu:2023yrb ; Li:2024joa ; Cassidy:2023lwd , searches for beyond-the-SM (BSM) heavy Higgs bosons Han:2022edd ; Bandyopadhyay:2020otm ; Han:2021udl ; Bandyopadhyay:2024plc ; Ouazghour:2023plc ; Jueid:2023qcf , investigations of dark matter Han:2020uak ; Capdevilla:2021fmj ; Han:2022ubw ; Cesarotti:2024rbh ; Asadi:2023csb ; Belfkir:2023vpo ; Jueid:2023zxx , constraints on lepton-universality violation Huang:2021nkl ; Huang:2021biu ; He:2024dwh , exploration of the muon $g-2$ anomaly Buttazzo:2020ibd ; Capdevilla:2020qel ; Yin:2020afe ; Capdevilla:2021rwo , and tests of a wide range of new physics scenarios Gu:2020ldn ; Costantini:2020stv ; Bandyopadhyay:2021pld ; Qian:2021ihf ; Bao:2022onq ; Ghosh:2023xbj ; Dermisek:2023tgq ; Altmannshofer:2023uci ; Liu:2023jta ; Sun:2023cuf ; Kwok:2023dck ; Cesarotti:2023sje ; Ema:2023buz ; Li:2023tbx ; Bhattacharya:2023beo ; Das:2024ekt .
@@
A muon collider is therefore both a discovery and precision machine. In particular, precision physics relies on both a clean experimental environment (no QCD in the initial state) and precise and reliable theoretical predictions. Thus, as has been the case for LEP, the calculation of electroweak (EW) corrections will be paramount for theoretical predictions in muon collisions. In fact, unlike other possible future electron-positron colliders, e.g. the Future Circular Collider (FCC-ee) FCC:2018evy ; Bernardi:2022hny and the Circular Electron-Positron Collider (CEPC) CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi ; An:2018dwb ; CEPCAcceleratorStudyGroup:2019myu ; CEPCPhysicsStudyGroup:2022uwl ; CEPCStudyGroup:2023quu , at a high-energy muon collider the scope and relevance of EW corrections will be much broader, especially for a 10 TeV (or higher) collision machine. [^1] Similar effects, although of smaller sizes due to the associated lower energies, are expected also for high-energy $e^{+}e^{-}$ colliders (see e.g. Ref. Kuhn:2007ca ) as CLIC, ILC and $\rm C^{3}$ ILC:2013jhg ; Behnke:2013lya ; ILCInternationalDevelopmentTeam:2022izu ; Linssen:2012hp ; Lebrun:2012hj ; CLIC:2016zwp ; Brunner:2022usy ; Bai:2021rdg ; Vernieri:2022fae . For such colliders, a more prominent effect originates from QED initial-state-radiation (ISR), which however in our set up would be automatically resummed within the PDF formalism. EW corrections at high energies can be very large, even of ${\cal O}(1)$ w.r.t. the leading-order (LO) prediction, and therefore are expected to be unavoidable in any phenomenological study, not only those regarding precision. The origin of such enhancements is the so-called EW Sudakov logarithms (EWSL), which involve logarithms of ratios of the form $Q/M_{W}$, where $Q$ is any of the scale of the process, like the energy of the collider $\sqrt{S}$, and $M_{W}$ is the $W$ boson mass.
@@
In this work, we precisely investigate this issue for the case of a muon collider at 3 and 10 TeV. We focus on direct production processes (also denoted in the literature as muon-muon annihilation), $\mu^{+}\mu^{-}\to F$, where the invariant mass of the final state is close to the energy of the collider and therefore VBF configurations are suppressed. In these configurations the only relevant PDFs are the ones of the (anti)muon in the (anti)muon, which accounts for effects from QED initial-state radiation (ISR). Moreover, our focus is not on precision physics but on large effects, as those expected from EWSL at high energies. We exploit the SM as a “test case”, but our conclusions are instructive also for a general BSM scenario. [^3] The case of SMEFT will be addressed in detail in an upcoming publication ElFaham:2024egs . All our calculations are performed via the MadGraph5_aMC@NLO framework Alwall:2014hca ; Frederix:2018nkq ; Pagani:2021vyk ; Bertone:2022ktl , in a completely automated approach.
@@
Then we approximate the resummation of EWSL via a simple exponentiation and, after additively matching them to the exact NLO EW, we compare this prediction with the exact NLO EW itself. We investigate when resummation is needed solely for precision studies and when it becomes essential to ensure sensible predictions and prevent negative cross sections.
@@ 2.1 The automation of NLO EW corrections
Given a physical observable, typically a cross section, the so-called NLO QCD and NLO EW corrections correspond to the exact ${\cal O}(\alpha_{s})$ and ${\cal O}(\alpha)$ corrections, respectively, to its LO prediction [^4] For the processes that we will consider in this paper, a single coupling combination contributes at LO. See Footnote 2 for the more general case.. Such corrections involve the calculation of one-loop amplitudes, their renormalisation, the regularisation of IR divergencies, and the combination of virtual as well as real-emission contributions in order to cancel them.
@@ 3.1 Calculation set-up
The two classes of processes entail completely different kinematics, especially in the bulk of the associated cross sections. Direct production is dominated by the phase-space region $s\simeq S$, where $\sqrt{s}$ is the total energy of the partonic process in its rest frame while $\sqrt{S}$ is the collider energy. Configurations with $s\neq S$ are induced by the emissions of photons, in particular the initial-state-radiation (ISR), which is taken into account in the collinear limit directly via PDF evolution of the (anti)muon in the (anti)muon $\Gamma_{\mu^{\pm}/\mu^{\pm}}$ or otherwise via the NLO EW corrections. VBF production is instead dominated by the phase-space region $\sqrt{s}\simeq m(F)\simeq\sum_{i\in F}m_{i}$, where $i$ is any particle that is part of the final state $F$ and $m(F)$ is the invariant mass of the final state $F$. In other words, The $\mu\mu^{-}/\nu_{\mu}\bar{\nu}_{\mu}$ additional pair in the final state carries away most of the energy of the colliding muons and it typically does it along the beam pipe axis. The hard process is in fact $VV\rightarrow F$, and the leading contributions, especially at very high energies, can be simulated in the so-called Effective-Vector-boson-Approximation (EVA) Dawson:1984gx ; Kane:1984bb , which has already been implemented in MadGraph5_aMC@NLO Ruiz:2021tdt . Instead of simulating $22\rightarrow n+$ 2 processes, where $n$ is the multiplicity of $F$, $2\rightarrow n$ matrix elements are sufficient ($VV\rightarrow F$) and the “$V$ in the muon” can be modeled similarly to what is done in the case of the photon in the lepton in the Weizsäcker–Williams approximation vonWeizsacker:1934nji ; Williams:1934ad . Not only, as already mentioned Sec. 2.1, these effects can also be resummed and taken into account in a PDF formalism, as shown in Refs. Han:2020uid ; Han:2021kes ; Garosi:2023bvq . [^9] It is interesting to note that since, at least for low multiplicities, $m(F)\simeq\sum_{i\in F}m_{i}\simeq M_{W}$, not only logarithms of the form $\log(s/M_{W}^{2})$ entering the PDFs are important but also power corrections of the form $(m(F)/M_{W})^{n}$ to the matrix elements cannot be neglected. However, this aspect is beyond the scope of this paper and it will not be investigated here.
@@
In this work we want to study the phenomenology of EW corrections at muon colliders for hard scattering processes at high energies, i.e., for the direct production mechanism.  [^10] The calculation of EW corrections for the direct production of multi-boson final states has also been performed in the Whizard framework Kilian:2007gr in Ref. Bredt:2022dmm . In particular, we consider final states with massive particles only. It is important to notice that unless a very high final-state multiplicity is chosen, in the SM for both 3 and 10 TeV collisions the two scales $m(F)$ and $S$ are very well separated, with $m(F)\ll S$. Thus, also when NLO EW corrections are taken into account, the two classes of processes (direct production and VBF) can be studied independently. In order to do so, unless differently specified, in all the results of the paper we apply the following cut on the invariant mass of the final state $F$:
@@
In order to mimic a realistic experimental set-up, since we consider only undecayed particles that are typically boosted, we require that they are both separated in angle among them and also with the beam pipe axis. In particular, for any particle $X$ and $Y$ that are part of the final state $F$ we require that
@@
where cuts are applied after the recombination of photons described before. These cuts should be considered as illustrative: for example, the $p_{T}(X)$ cut may be larger depending on the collider energy. Our overall conclusions are not affected by the specific values employed.
@@ 3.2.2 EWSL and their approximate resummation
The r.h.s. of Eq. (3.20) says that if $\sigma_{\rm EXP_{EW}}$ is expanded in powers of $\alpha$ the NLO EW prediction is captured exactly, while beyond ${\cal O}(\alpha)$ the resummed tower of EWSL of order $\alpha^{n}\log^{k}(s/M_{W}^{2})$ with $n>1$ and $k=2n,2n-1$ is approximated via simple exponentiation. We stress again that we do not claim we are doing NLL resummation of EWSL. We instead want to study when and if this procedure is necessary, by comparing $\delta_{\rm NLO_{EW}}$ with the relative corrections induced by $\sigma_{\rm EXP_{EW}}$, namely
@@ 3.3 List of aspects investigated in this work
    When NLO EW corrections reach, or even worse surpass, the relative size of $-100\%$ of the LO, resummation of EWSL does not concern precision but the physical sense of predictions. We want to investigate when we should expect such situations at 3 and 10 TeV collisions.
@@ 4 Numerical results
In this section we will also start to describe features of the EW corrections that are distinctive for direct production processes at high-energy lepton colliders and quite different from the case of a high-energy hadronic machine such as LHC or FCC-hh (see, e.g., Refs. Mangano:2016jyj ; Azzi:2019yne for overviews of EW corrections at high-energy hadronic machines). We remind the reader that all calculations have been performed for SM processes, which are the only ones that can be calculated at $\rm NLO_{EW}$ accuracy with automated tools. However, many of the conclusions from this study, especially those stating the limitations of some approaches, can be clearly generalised also to BSM scenarios. The SM has to be regarded only as a test case in this respect.
@@ 4.1.1 Importance of ${\rm SDK_{weak}}$ versus ${\rm SDK_{0}}$
Using the same colour code of Ref. Pagani:2021vyk , in the main panel we show LO (blue), $\rm NLO_{EW}$ (orange), ${\rm SDK_{weak}}$ (red), and ${\rm SDK_{0}}$ (green) predictions. [^16] The rigorous definition of these quantities can be found in Sec. 3.2. Whenever a distribution has a negative cross section, the corresponding histogram is plotted as a dotted line. In the first inset, we plot the relative corrections induced by such approximations w.r.t. the LO predictions, i.e., the quantities $\delta_{X}$ defined in Eqs. (3.10), (3.14) and more in general (3.26). In such inset it is possible to appreciate the size and the sign of EW corrections, either calculated at exact $\rm NLO_{EW}$ accuracy or via Sudakov approximation(s). Then, in the second inset, we plot the difference $\delta_{X}-\delta_{\rm NLO_{EW}}$ for the two cases $X={\rm SDK_{weak}}$ and $X={\rm SDK_{0}}$ in order to test their accuracy. Rather than the minimisation of this quantity, the validity of the Sudakov approximation consists in having a small constant difference ($|\delta_{X}-\delta_{\rm NLO_{EW}}|\simeq\alpha$) over the full spectrum, i.e. a horizontal line, see also Eqs. (3.18) and (3.19) and text around it. Indeed, ${\cal O}(\alpha)$ contributions are expected to be present, while non-horizontal lines indicate an (at least) logarithmic-enhanced contribution that is not captured. Such contribution may accidentally compensate the ${\cal O}(\alpha)$ constant term and lead to $\delta_{X}-\delta_{\rm NLO_{EW}}\simeq 0$ for particular phase-space regions, but this is not to be regarded as an indication of the validity of the approximation. That said, a large constant difference ($|\delta_{X}-\delta_{\rm NLO_{EW}}|\gg\alpha$), however, also points to logarithms that are not correctly captured. In particular, at a high-energy lepton collider, the direct-production processes studied in this work and characterised by $\sqrt{s}\simeq\sqrt{S}$ may show such effects induced by missing large double logarithms (see Eq. (2.3)) of the form $L(s,M_{W}^{2})$, [^17] It is easy to see in Sec. 4.1 of Ref. Pagani:2021vyk that, at variance with the ${\rm SDK_{0}}$ scheme, the ${\rm SDK_{weak}}$ scheme implies the substitution $C_{\rm EW}\to C_{\rm EW}-Q^{2}$, where $Q$ is the charge of the particle considered, in the prefactor of the double-logarithm of the form $L(s,M_{W}^{2})$ entering the formula of $\delta_{\rm DL}$ in Eq. (3.14). Similar effects are also present in the single logarithms and lead to large constants when a fixed $s$ is considered. which are therefore large constants for the full phase-space.
@@
Figure 5: Same as Fig. 1, for the $p_{T}(H)$ distribution in $\mu^{+}\mu^{-}\to ZH$.
Finally, in Fig. 5 we show the transverse-momentum distribution for the Higgs boson in $\mu^{+}\mu^{-}\to ZH$ production, $p_{T}(H)$. Similarly to the case of $t\bar{t}$ production in Fig. 1, we observe a much less flat LO prediction w.r.t. the case of $WW$ and $ZZ$ production. As in the case of $t\bar{t}$ production, and unlike $WW$ and $ZZ$, the tree-level amplitude features only an $s$-channel diagram. As in $t\bar{t}$ production, and even more, $\delta_{\rm NLO_{EW}}$ also is quite flat over the full spectrum. We also observe that the difference $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$ is larger, of the order of $10\%$ both for 3 and 10 TeV collisions. Instead, in the case of $\delta_{{\rm SDK_{0}}}-\delta_{\rm NLO_{EW}}$, such difference is of the order of $15\%$ at 3 TeV and $20\%$ at 10 TeV. Thus, the accuracy of $\delta_{{\rm SDK_{0}}}$ is not only worse but also energy-dependent. To the best of our understanding, the $\mu^{+}\mu^{-}\to ZH$ has large non-logarithmic-enhanced contributions at ${\cal O}(\alpha)$ and on top of that the $\delta_{{\rm SDK_{0}}}$ wrongly approximates the double logarithms of the form $L(s,M_{W}^{2})\simeq L(S,M_{W}^{2})$ from the photon exchange among the muons in the initial state. Notice that the difference $\delta_{{\rm SDK_{weak}}}-\delta_{{\rm SDK_{0}}}$ is the same observed, both at 3 and 10 TeV, for $\mu^{+}\mu^{-}\to ZZ$ production in Fig. 4, for which the same argument was presented.
We have also calculated and analysed the $m(ZH)$ distribution. The situation is similar to the one observed for the $m(ZZ)$ distribution, but the $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$ remains constant at the order of $10\%$ as discussed for the $p_{T}(H)$ distribution. Some of the features discussed in this section have also been observed in Ref. Bredt:2022dmm , where for the $ZZ$ and $ZH$ final state one can also find the analytical expression of the corresponding EWSL.